# The RAG Paradox: A Black-Box Attack Exploiting Unintentional Vulnerabilities in Retrieval-Augmented Generation Systems

### v2

@@ Abstract
Chanwoo Choi1 Jinsoo Kim1 Sukmin Cho2 Soyeong Jeong2 Buru Chang3,thanks: Corresponding author. 1Sogang University  2KAIST  3Korea University {cw316,jinsoolve}@sogang.ac.kr, smcho@casys.kaist.ac.kr starsuzi@kaist.ac.kr, buru_chang@korea.ac.kr
@@ 3.1 Threat Model
Attacker’s goal. The attacker aims to prevent the RAG system from generating the correct answer for a set of target queries. To achieve this, the attacker pursues three key objectives. First, the attacker crafts poisoned documents to be highly retrievable. Second, the retrieved documents are designed to interfere with the answer generation process, causing the system to produce incorrect or misleading responses. Third, the attacker ensures that the poisoned documents appear natural and coherent, so that even when presented to users as sources, they do not raise suspicion about the generated responses. This combination of goals enables a highly effective and difficult-to-detect black-box attack against real-world RAG systems.
Attacker’s capabilities. We assume an attacker with no internal access to the target system. However, based on the RAG paradox, the attacker can query the RAG system to obtain the retrieved documents and their disclosed sources. By analyzing these documents, the attacker can infer the retriever’s preferred phrasing. Additionally, the attacker can identify external platforms referenced by the system, such as Wikipedia, Reddit, and LinkedIn, and upload content to these platforms.
@@ 3.2 Our Attack Scenario
Poisoned Document Publishing. We publish the poisoned documents on external platforms—such as Wikipedia, Reddit, and LinkedIn—that were previously identified in the RAG system’s responses. Once indexed, these documents can be retrieved by the system, providing an entry point for external attackers to manipulate its behavior.
@@ 3.3 Poisoned Document Generation with PARADOX
Based on these considerations, our poisoning method uses the Llama-3.1-8B-Instruct model to generate poisoned documents in the following steps. Appendix §$\S$§A provides details of our method, including the prompts used.
@@ 3.3.2 Document Generation
Dataset Method NES: ↑ (better) Doc Selection Rate NDCG@5
@@
Dataset Method Accuracy: ↓ (better) & ASR: ↑ (better) Doc Selection Rate & NDCG@5
@@
Ours (w/o E2D) 33.33 | 65.41 34.43 | 65.09 36.08 | 63.84 33.02 | 66.90 3.98 | 0.83 3.80 | 0.80 4.74 | 0.96 4.57 | 0.93
Table 3: Ablation test results using Llama-3.1-8B-Instruct.(*) indicates ($p<0.05$), (**) indicates ($p<0.01$),
@@ 4 Experiments
To validate the effectiveness and feasibility of our realistic attack scenario, we conduct offline experiments using datasets and generators commonly used in RAG research. We further perform a limited number of carefully controlled online experiments, conducted solely for research purposes to ensure safety and ethical compliance, targeting commercial RAG systems. These experiments confirm that our attack method is effective in real-world deployment settings. The details of our experiments are provided in Appendix §$\S$§B.
@@ 4.2 Experimental Results
Ablation test. We conducted an ablation study to verify the effectiveness of Retriever Selection Analysis. As shown in Table 3, incorporating Retriever Selection Analysis consistently resulted in lower accuracy, while achieving higher ASR and document selection rates across all retrievers and datasets. These results confirm that Retriever Selection Analysis enhances the effectiveness of the attack by increasing the retrievability of poisoned documents. Notably, the effect is most pronounced when BM25 is used as the retriever, which we attribute to its ability to effectively identify and emphasize key phrases that influence BM25’s sparse matching mechanism. Statistical analysis further supports this, showing significant increases in the average number of retrieved poisoned documents, with p-values mostly below 0.01. Overall, these results show that Retriever Selection Analysis is important for making poisoned documents more likely to be retrieved and for causing bigger performance drops in the system.
@@
Fictional Indv. 75% 100$\rightarrow$40100% 100$\rightarrow$30%
Rare Species 25% 100$\rightarrow$75100% 100$\rightarrow$30%
Everyday Questions 10% 100$\rightarrow$9050% 100$\rightarrow$50%
Product Review 10% 100$\rightarrow$9070% 100$\rightarrow$30%
@@
We prepare five QA pairs per target each for Fictional Individuals and Rare Species, and ten pairs each for Everyday Questions and Product Review. Next, we verify that ChatGPT and Perplexity generate accurate responses using retrieved documents for the QA pairs. Then we apply our attack method and upload the poisoned documents to external sources. Finally, we evaluate the document selection rate (SR) of poisoned documents per query. SR indicates cases where poisoned documents appear as sources or citations in the response. We assess this metric in both RAG systems and analyze the resulting performance changes. Appendix §$\S$§D provides details on QA pairs.
@@ B.1 Implementation Details
[…]
@@ B.2 Template
B.3 Details of NES
@@
Dataset Method Accuracy: ↓ (better) ASR: ↑ (better)
@@
Dataset Method Accuracy: ↓ (better) ASR: ↑ (better)
@@ D.2.4 Product Review
Figure 5: Prompts used for Retriever Behavior Inference and Document Generation.

### v3

@@ Abstract
Chanwoo Choi1 Jinsoo Kim2 Sukmin Cho3 Soyeong Jeong3 Buru Chang1,thanks: Corresponding author. 1Korea University  2Sogang University  3KAIST {ccw316,buru_chang}@korea.ac.kr, jinsoolve@sogang.ac.kr smcho@casys.kaist.ac.kr, starsuzi@kaist.ac.kr
@@ 3.1 Threat Model
Attacker’s goal. The attacker aims to prevent the RAG system from generating the correct answer for a set of target queries. In particular, we consider RAG systems that retrieve documents from public sources as primary targets. To achieve this, the attacker pursues three key objectives. First, the attacker crafts poisoned documents to be highly retrievable. Second, the retrieved documents are designed to interfere with the answer generation process, causing the system to produce incorrect or misleading responses. Third, the attacker ensures that the poisoned documents appear natural and coherent, so that even when presented to users as sources, they do not raise suspicion about the generated responses. This combination of goals enables a highly effective and difficult-to-detect black-box attack against real-world RAG systems.
Attacker’s capabilities. We assume an attacker with no internal access to the target system. However, based on the RAG paradox, the attacker can query the RAG system to obtain the retrieved documents and their disclosed sources. By analyzing these documents, the attacker can infer the retriever’s preferred phrasing. Additionally, the attacker can identify external platforms referenced by the system, such as Wikipedia, Reddit, and LinkedIn, and upload content to these platforms. This capability is limited to posting documents on the identified platforms, without extending to any direct control over how the system subsequently indexes or integrates such content.
@@ 3.2 Our Attack Scenario
Poisoned Document Publishing. We publish the poisoned documents on external platforms—such as Wikipedia, Reddit, and LinkedIn—that were previously identified in the RAG system’s responses. Once the system indexes the uploaded documents and they become searchable, these documents can be retrieved by the system, providing an entry point for external attackers to manipulate its behavior.
@@ 3.3 Poisoned Document Generation with PARADOX
[…]
@@ 3.3.2 Document Generation
Dataset Method NES (↑ better) Doc Selection Rate NDCG@5
@@
Dataset Method Accuracy (↓ better) & ASR (↑ better) Doc Selection Rate & NDCG@5
@@
Ours (w/o RPA) 33.33 | 65.41 34.43 | 65.09 36.08 | 63.84 33.02 | 66.90 3.98 | 0.83 3.80 | 0.80 4.74 | 0.96 4.57 | 0.93
Table 3: Ablation test results using Llama-3.1-8B-Instruct.(*) indicates ($p<0.05$), (**) indicates ($p<0.01$). RPA refers to Retriever Preference Analysis.
Dataset Method Reranking: Accuracy (↓ better) Confidence Reasoning: Accuracy (↓ better)
BM25 Contriever BM25 Contriever
NQ Clean 37.12 40.58 40.00 47.00
PoisonedRAG-$B$ 8.92 (-76%) 9.64 (-76%) 22.00 (-45%) 19.00 (-60%)
Vec2Text 35.15 (-5%) 31.75 (-22%) 43.00 (+8%) 36.00 (-23%)
HotFlip 10.25 (-72%) 8.34 (-79%) 20.00 (-50%) 22.00 (-53%)
Ours 5.04 (-86%) 6.48 (-84%) 15.00 (-62%) 15.00 (-68%)
HotpotQA Clean 38.10 35.58 34.00 34.00
PoisonedRAG-$B$ 6.22 (-83%) 6.28 (-82%) 15.00 (-56%) 14.00 (-59%)
Vec2Text 36.15 (-5%) 22.40 (-37%) 25.00 (-26%) 21.00 (-38%)
HotFlip 6.69 (-82%) 5.52 (-84%) 11.00 (-68%) 11.00 (-68%)
Ours 2.81 (-93%) 1.93 (-95%) 10.00 (-71%) 8.00 (-76%)
MedQA Clean 45.36 46.31 35.00 34.00
PoisonedRAG-$B$ 51.73 (+14%) 51.10 (+10%) 46.00 (+31%) 47.00 (+38%)
Vec2Text 46.23 (+2%) 44.10 (-5%) 35.00 (0%) 34.00 (0%)
HotFlip 47.96 (+6%) 47.48 (+3%) 44.00 (+26%) 49.00 (+44%)
Ours 29.87 (-34%) 32.86 (-29%) 27.00 (-23%) 30.00 (-12%)
Table 4: Attack effectiveness under two defense methods: Reranking Yoon et al. (2024) and Confidence Reasoning Huang et al. (2025).
@@ 4 Experiments
[…]
@@ 4.2 Experimental Results
Ablation test. We conduct an ablation study to verify the effectiveness of Retriever Preference Analysis. As shown in Table 3, incorporating Retriever Preference Analysis consistently resulted in lower accuracy, while achieving higher ASR and document selection rates across all retrievers and datasets. These results confirm that Retriever Preference Analysis enhances the effectiveness of the attack by increasing the retrievability of poisoned documents. Notably, the effect is most pronounced when BM25 is used as the retriever, which we attribute to its ability to effectively identify and emphasize key phrases that influence BM25’s sparse matching mechanism. Statistical analysis further supports this, showing significant increases in the average number of retrieved poisoned documents, with p-values mostly below 0.01, and we further provide additional quantitative analysis on Retriever Preference Analysis in Appendix B.3.
Overall, these results show that Retriever Preference Analysis is important for making poisoned documents more likely to be retrieved and for causing bigger performance drops in the system.
Attack effectiveness under defenses. We further evaluate the proposed attack within defense-integrated RAG systems by applying two representative defenses: re-ranking Yoon et al. (2024) and confidence reasoning Huang et al. (2025). All experiments use Llama-3.1-8B-Instruct with BM25 as the sparse retriever and Contriever as the dense retriever. For re-ranking, we retrieve the top-50 documents per query and apply tournament-style re-ranking with ListT5-base Yoon et al. (2024); we then assess attack effectiveness on the top-5 re-ranked documents across the full query set. Re-ranking aims to defend by demoting poisoned documents that are unhelpful to the generator, thereby reducing their influence on final generations.
For confidence reasoning, we adopt rule-based confidence reasoning Huang et al. (2025) evaluated on 100 randomly selected queries. Confidence reasoning defends by detecting when retrieved documents do not meaningfully improve generation quality and by omitting such documents from the generation process.
As shown in Table 4, most existing attacks remain vulnerable even after re-ranking, whereas our method consistently produces the largest performance drop. On MedQA, some existing attacks even increase performance relative to the clean corpus, yet our method still degrades system performance. While confidence reasoning partially mitigates the attack impact, our attack continues to induce the largest drop. These findings suggest that the poisoned documents generated by our method are (i) still ranked as relevant by the reranker and (ii) assessed by the confidence filter as sufficiently helpful for generation, allowing them to survive both defenses. Overall, the attack remains highly effective across diverse retrieval settings and defensive mechanisms, demonstrating its robustness and practical impact.
@@
[…]
@@
[…]
@@ 5 Conclusion
Acknowledgment
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation(IITP)-ICT Creative Consilience Program grant funded by the Korea government(MSIT)(IITP-2025-RS-2020-II201819) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2024-00350430, RS-2025-24533089).
@@ B.1 Implementation Details
  • HotFlip. We adopt the white-box setting from PoisonedRAG, using HotFlip as the optimization method. We set $\texttt{num\_iter}=10$ and $\texttt{num\_cand}=20$ for HotFlip.
@@ B.2 Template
Dataset Retriever Mean Difference Standard Error 95% Confidence Interval
NQ BM25 +0.6787 0.0182 (0.6429, 0.7144)
Contriever +0.2241 0.0171 (0.1907, 0.2575)
ANCE +0.1396 0.0124 (0.1153, 0.1639)
BGE +0.0194 0.0105 (-0.0012, 0.0399)
HotpotQA BM25 +0.2493 0.0087 (0.2322, 0.2664)
Contriever +0.0105 0.0035 (0.0036, 0.0174)
ANCE +0.0608 0.0079 (0.0453, 0.0762)
BGE +0.0367 0.0081 (0.0208, 0.0527)
MedQA BM25 +0.2453 0.0244 (0.1973, 0.2932)
Contriever +0.1077 0.0298 (0.0492, 0.1662)
ANCE +0.0432 0.0127 (0.0183, 0.0681)
BGE +0.1297 0.0188 (0.0928, 0.1667)
Table 8: Mean difference, standard error, and 95% confidence intervals for different retrievers across datasets.
B.3 Quantitative Analysis of Retriever Preference Analysis
We conduct a quantitative analysis to evaluate the effectiveness of Retriever Preference Analysis. Using paired t-tests, we confirm that in most cases the improvements are statistically significant, while in some conditions the significance is limited. Table 8 shows that the proposed method generally increases the frequency of poisoned documents being retrieved, with the largest effects observed in sparse retrievers. In contrast, the improvements in dense retrievers are relatively smaller, yet still consistent and reliable. Table 9 shows a similar trend, where Retriever Preference Analysis yields the most pronounced improvements in ASR for sparse retrievers. For dense retrievers, the magnitude of improvement is more limited, and some results are not statistically significant, yet an overall consistent pattern of gains is still observed. These findings demonstrate that Retriever Preference Analysis systematically enhances attack performance across different retriever types, with the most substantial effects observed in sparse retrievers.
Dataset Retriever Mean Difference Standard Error 95% Confidence Interval
NQ BM25 +4.16 0.53 (3.12, 5.19)
Contriever +1.36 0.51 (0.36, 2.35)
ANCE +0.47 0.41 (-0.33, 1.27)
BGE +0.58 0.41 (-0.22, 1.38)
HotpotQA BM25 +0.50 0.24 (0.02, 0.98)
Contriever +0.38 0.20 (-0.01, 0.77)
ANCE +0.42 0.24 (-0.06, 0.89)
BGE +0.23 0.25 (-0.27, 0.73)
MedQA BM25 +3.07 1.55 (0.03, 6.10)
Contriever +2.28 1.63 (-0.91, 5.47)
ANCE +0.71 1.61 (-2.46, 3.87)
BGE +2.20 1.58 (-0.91, 5.31)
Table 9: Mean difference, standard error, and 95% confidence intervals for different retrievers across datasets.
Figure 4: Additional NES Evaluation
B.4 Details of NES
@@
Additionally, we provide Figures 7, 8, 9, 10, and 11, which illustrate document examples corresponding to NES scores from 1 to 5 along with the evaluations made by the LLM evaluator. Each figure includes poisoned documents generated by four different attack methods, thereby demonstrating how the evaluator interprets these documents and assigns the corresponding NES scores. These examples highlight the concrete evaluation process and provide clearer evidence for the consistency and validity of the assigned scores.
@@
Dataset Method Accuracy (↓ better) ASR (↑ better)
@@
Dataset Method Accuracy (↓ better) ASR (↑ better)
@@ D.2.4 Product Review
Figure 5: Prompts used for Retriever Preference Analysis and Document Generation.
@@
Figure 7: Examples of Document and LLM Evaluation when NES = 1
Figure 8: Examples of Document and LLM Evaluation when NES = 2
Figure 9: Examples of Document and LLM Evaluation when NES = 3
Figure 10: Examples of Document and LLM Evaluation when NES = 4
Figure 11: Examples of Document and LLM Evaluation when NES = 5